in a mixture of gold and silver , silver is 38.5% of total weight.its density of 14.6gram per ml.what is the molar concentration of silver in it mixture?
1
Expert's answer
2017-06-27T05:22:02-0400
Answer on Question #68936, Chemistry / General Chemistry
in a mixture of gold and silver , silver is 38.5% of total weight.its density of 14.6gram per ml.what is the molar concentration of silver in it mixture?
Answer
n(Ag)=m/M=38.5/108=0.356Mol
V=m/₰=100g/14.6g/Mol=6.849Ml
Cm(Ag)=0.356Mol*1000Ml/6.849Ml=51.98Mol/L